OpenStudy (anonymous):

Solve the equation: (sin^-1 x) - cos ^-1(sqrt(3)/3) = 0 What do I do with sin after I convert cos?

OpenStudy (anonymous):

use \[\cos^{-1} A+\cos^{-1} B=\cos^{-1} (AB+\sqrt{1-B^2}\sqrt{1-A^2})\]

OpenStudy (anonymous):

I don't understand, I keep getting strange answers

OpenStudy (anonymous):

\[\sin^{-1} (x) -\cos^{-1} (\sqrt{3}/3) =0\] \[\sin^{-1} (x) =\cos^{-1} (\sqrt{3}/3) \] \[ x =\sin (\cos^{-1} (\sqrt{3}/3) )\] ?